solve log(x+9)-logx=3
Answer by stanbon(75887)

solve log(x+9)-logx=3
log[(x+9)/x] = 3
(x+9)/x = 10^3
x+9 = 1000x
999x = 9
x = 1/333
==================
Cheers,
Stan H.